How fast is the 5.00 kg pumpkin from the previous question moving when it is 1.00 m above the ground (before it hits)?
1
Expert's answer
2020-12-10T11:03:46-0500

Assume that the pumpkin falls from a height of "H=10 m" without initial velocity.

So, its speed at a distance of 1 m above the ground


"mgH_0=mgh+mv^2/2\to g(H_0-h)=v^2/2\to"


"v=\sqrt{2g(H_0-h)}=\sqrt{2\cdot9.8\cdot (10-1)}=13.3(m)"